OpenStudy (anonymous):

Simplify the expression .. sqrt45n^5 I'll put the equation in the comment thingy..

OpenStudy (anonymous):

\[\sqrt{45n ^{5}}\]

OpenStudy (anonymous):

\[3n^2\sqrt{5n}\]

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

and I'll brb

OpenStudy (anonymous):

\[\sqrt{9*5*n ^{4}*n}=3n^2\sqrt{5*n}\]

OpenStudy (anonymous):

firstli factorise inside the sqrt, then separate them autside.. is it ok ?

OpenStudy (anonymous):

Thank you